# The $S_{n}$-equivariant Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$

### v1

@@ 7.1. Calculation of $\mathsf{b}_{1,r}^{\mathrm{nrt}}$
Their specialisations to symmetric functions are denoted as $\mathsf{a}_{0}^{\prime\prime}$ and $\dot{\mathsf{a}}_{0}$, respectively. Therefore, to determine $\mathsf{e}(\mathsf{Dih}_{r})\circ_{S_{2}}\mathsf{e}(\mathsf{Cat})$, it suffices to plug in formulas of $\mathsf{a}_{0}^{\prime\prime}$ and $\dot{\mathsf{a}}_{0}$ to Theorem 6.4. To do this, we differentiate the following formula of Getzler [Get95a]:
@@
The term $\mathscr{B}_{0}$ specialises to $(r+1)\sum_{n\geq 1}\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}})$, which is determined in [Get98] and [Pet12] in terms of $\mathsf{a}_{0}^{\prime\prime}$, $\dot{\mathsf{a}}_{0}$, and $\mathsf{a}_{1}=\sum_{n=1}^{\infty}\chi^{S_{n}}(\mathcal{M}_{1,n})$. A formula for $\mathsf{a}_{1}$ is derived by Getzler in [Get99, (5.5)]. Plugging these formulas into Theorem 6.4, we obtain Table 5 for degree 222 maps and Table 6 for degree 333 maps.
7.2. Contribution of rational tails
For small values of $n$ and $d$, the $S_{n}$-equivariant Euler characteristic of the Kontsevich space ¯$\overline{\mathcal{M}}_{0,n}(\mathbb{P}^{r},d)$ can be determined explicitly via torus localisation. Performing the plethysm
$\mathsf{b}_{1,r}^{\mathrm{nrt}}\circ\left(p_{1}+\frac{\mathsf{b}_{0,r}^{\prime}}{r+1}\right)$
produces Tables 1, 2, and 3. The sample calculations reflect geometric properties of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. For instance, Table 7 compares the contributions from the two terms
(7.5) $\chi^{S_{n}}(\overline{\mathcal{M}}^{\mathrm{nrt}}_{1,n}(\mathbb{P}^{r},d))\text{ and }\mathsf{b}_{1,r}^{\mathrm{rt}}(n,d):=\left[\operatorname{ch}(\overline{\mathcal{M}}^{\mathrm{nrt}}_{1}(\mathbb{P}^{r},0))\circ(p_{1}+\mathsf{b}_{0,r}^{\prime})\right]_{(n,d)}$
[…]
Regarding the lack of smoothness of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$, it may be of interest to calculate the $S_{n}$-equivariant Euler characteristics of all smoothable genus one stable maps. We hope to revisit this in relation to the Vakil–Zinger desingularisation of the mapping space [VZ08], [RSPW19].
@@ 0 $4\binom{r+1}{2}$
Table 1. The $S_{n}$-equivariant topological Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},1)$, for $n\leq 3$.
@@ 0 $24\binom{r+1}{3}+17\binom{r+1}{2}$
$+\left[144\binom{r+1}{3}+40\binom{r+1}{2}+6\binom{r+1}{1}\right]s_{1,1,1}$
Table 2. The $S_{n}$-equivariant topological Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},2)$, for $n\leq 3$.
@@ 0 $216\binom{r+1}{4}+247\binom{r+1}{3}+55\binom{r+1}{2}$
Table 3. The $S_{n}$-equivariant topological Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},3)$, for $n\leq 2$.
$d$ $\mathsf{e}\left(\overline{\mathcal{M}}_{1,0}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}\right)$
222 $\binom{r+1}{2}$
333 $\binom{r+1}{3}+\binom{r+1}{2}$
444 $3\binom{r+1}{4}+6\binom{r+1}{3}+3\binom{r+1}{2}$
555 $12\binom{r+1}{5}+24\binom{r+1}{4}+15\binom{r+1}{3}+3\binom{r+1}{2}$
666 $60\binom{r+1}{6}+150\binom{r+1}{5}+136\binom{r+1}{4}+53\binom{r+1}{3}+8\binom{r+1}{2}$
777 $360\binom{r+1}{7}+1080\binom{r+1}{6}+1200\binom{r+1}{5}+600\binom{r+1}{4}+129\binom{r+1}{3}+9\binom{r+1}{2}$
888 $2520\binom{r+1}{8}+8820\binom{r+1}{7}+11970\binom{r+1}{6}+7905\binom{r+1}{5}+2607\binom{r+1}{4}+393\binom{r+1}{3}+21\binom{r+1}{2}$
Table 4. The number of $\mathbb{C}^{*}$-fixed points of ¯$\overline{\mathcal{M}}_{1,0}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$ as a polynomial in $r$.
$n$ $\chi^{S_{n}}\left(\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},2)\right)$
111 $2\binom{r+1}{2}s_{1}$
222 $3\binom{r+1}{2}s_{2}+\binom{r+1}{2}s_{1,1}$
333 $4\binom{r+1}{2}s_{3}+2\binom{r+1}{2}s_{2,1}$
444 $7\binom{r+1}{2}s_{4}+3\binom{r+1}{2}s_{3,1}+\binom{r+1}{2}s_{2,2}$
Table 5. The first 444 coefficients of $q^{2}$ in $\mathsf{b}_{1,r}^{\mathrm{nrt}}$.
$n$ $\chi^{S_{n}}\left(\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},3)\right)$
111 $\left[3\binom{r+1}{3}+2\binom{r+1}{2}\right]s_{1}$
222 $\left[6\binom{r+1}{3}+3\binom{r+1}{2}\right]s_{2}+\left[3\binom{r+1}{3}+\binom{r+1}{2}\right]s_{1,1}$
333 $\left[13\binom{r+1}{3}+6\binom{r+1}{2}\right]s_{3}+\left[8\binom{r+1}{3}+2\binom{r+1}{2}\right]s_{2,1}+\binom{r+1}{3}s_{1,1,1}$
444 $\left[27\binom{r+1}{3}+11\binom{r+1}{2}\right]s_{4}+\left[24\binom{r+1}{3}+7\binom{r+1}{2}\right]s_{3,1}+\left[3\binom{r+1}{3}-\binom{r+1}{2}\right]s_{2,2}+3\binom{r+1}{3}s_{2,1,1}$
Table 6. The first 444 coefficients of $q^{3}$ in $\mathsf{b}_{1,r}^{\mathrm{nrt}}$.
$n$ $\chi^{S_{n}}\left(\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},2)\right)$ $\mathsf{b}_{1,r}^{\mathrm{rt}}(n,2)$
0 0 $24\binom{r+1}{3}+17\binom{r+1}{2}$
111 $2\binom{r+1}{2}s_{1}$ $\left[108\binom{r+1}{3}+56\binom{r+1}{2}\right]s_{1}$
222 $3\binom{r+1}{2}s_{2}+\binom{r+1}{2}s_{1,1}$ $\left[339\binom{r+1}{3}+168\binom{r+1}{2}+6\binom{r+1}{1}\right]s_{2}+\left[168\binom{r+1}{3}+64\binom{r+1}{2}+6\binom{r+1}{1}\right]s_{1,1}$
3 $4\binom{r+1}{2}s_{3}+2\binom{r+1}{2}s_{2,1}$
$\left[1176\binom{r+1}{3}+494\binom{r+1}{2}+6\binom{r+1}{1}\right]s_{3}+\left[996\binom{r+1}{3}+394\binom{r+1}{2}+12\binom{r+1}{1}\right]s_{2,1}$
$+\left[144\binom{r+1}{3}+40\binom{r+1}{2}+6\binom{r+1}{1}\right]s_{1,1,1}$
Table 7. Comparison between maps without rational tails and maps which contract the genus one core. The term $\mathsf{b}_{1,r}(n,2)$, defined in (7.5), is the contribution from maps which contract the core.

### v2

@@ 7. Calculations
Their specialisations to symmetric functions are denoted as $\mathsf{a}_{0}^{\prime\prime}$ and $\dot{\mathsf{a}}_{0}$, respectively. Therefore, to determine $\mathsf{e}(\mathsf{Dih}_{r})\circ_{S_{2}}\mathsf{e}(\mathsf{Cat})$, it suffices to plug in formulas for $\mathsf{a}_{0}^{\prime\prime}$ and $\dot{\mathsf{a}}_{0}$ into Theorem 6.4. To do this, we differentiate the following formula of Getzler [22]:
@@
The term $\mathscr{B}_{0}$ specialises to $(r+1)\sum_{n\geq 1}\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}})$, which is determined in [23] and [54] in terms of $\mathsf{a}_{0}^{\prime\prime}$, $\dot{\mathsf{a}}_{0}$, and $\mathsf{a}_{1}=\sum_{n=1}^{\infty}\chi^{S_{n}}(\mathcal{M}_{1,n})$. A formula for $\mathsf{a}_{1}$ is derived by Getzler in [24, (5.5)]. Plugging these formulas into Theorem 6.4 and taking the plethysm with $p_{1}+\mathsf{b}_{0,r}^{\prime}/(r+1)$, we obtain Tables 1, 2, and 3. Several additional tables, as well as a discussion of the contribution from excess components to $\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d))$ can be found in the first arXiv version of the present article.
@@ 33 $86\binom{r+1}{2}s_{3}+56\binom{r+1}{2}s_{2,1}+4\binom{r+1}{2}s_{1,1,1}$
Table 1. The $S_{n}$-equivariant topological Euler characteristic of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},1)$, for $n\leq 3$.
@@ 11 $\left[108\binom{r+1}{3}+58\binom{r+1}{2}\right]s_{1}$
$\left[144\binom{r+1}{3}+40\binom{r+1}{2}+6\binom{r+1}{1}\right]s_{1,1,1}$
Table 2. The $S_{n}$-equivariant topological Euler characteristic of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},2)$, for $n\leq 3$.
@@ 11 $\left[1300\binom{r+1}{4}+1365\binom{r+1}{3}+260\binom{r+1}{2}\right]s_{1}$
Table 3. The $S_{n}$-equivariant topological Euler characteristic of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},3)$, for $n\leq 2$.